OpenStudy (anonymous):

What is the solution of x + y = 2 and 2x + 5y = 16? a.) x=1, y=4 b.)x=4, y=-2 c.)x=-2, y=-2 d.)x=-2, y=4 HELP!!!!!!

OpenStudy (imstuck):

To do this, you can either use substitution or the addition method. Here it will probably be easier to use substitution. Look at the first equation: x + y = 2. If you solve this equation for y, you now know y's value in terms of x. y = 2 - x. Now that you "know" what y is, substitute it into your second equation.\[y=2-x-->2x+5(2-x)=16\]Now you have an equation that only has one variable in it instead of 2. Distribute the 5 into your (2 - x) to get 10 - 5x. Now your equation looks like this:\[2x+10-5x=16\]Now, do the math! 2x - 5x = -3x. -3x + 10 = 16. Subtract the 10 from both sides and you have -3x = 6, and solving for x gives you -2. NOW you know x and you can find y. Substitute -2 in for x in either of your equations (I would pick the first one because it's simpler to solve). x + y = 2 yields -2 + y = 2. Now solve for y. Add 2 to both sides and y = 4. See that?